OpenStudy (anonymous):

What is the probability of rolling 5 or 6 on a number cube? 45%, 30%, 25%,33

OpenStudy (campbell_st):

P(5) = 1/6 P(5 or 6) = 2/6 = 1/3 1/3 = 33% that is providing there is one 5 and one 6 on the cube

OpenStudy (anonymous):

thank you

OpenStudy (campbell_st):

hope it helps